# Learning policies for resource allocation in business processes

### v2

@@ Abstract
Efficient allocation of resources to activities is pivotal in executing business processes but remains challenging. While resource allocation methodologies are well-established in domains like manufacturing, their application within business process management remains limited. Existing methods often do not scale well to large processes with numerous activities or optimize across multiple cases. This paper aims to address this gap by proposing two learning-based methods for resource allocation in business processes. The first method leverages Deep Reinforcement Learning (DRL) to learn near-optimal policies by taking action in the business process. The second method is a score-based value function approximation approach, which learns the weights of a set of curated features to prioritize resource assignments. To evaluate the proposed approaches, we first designed six distinct business processes with archetypal process flows and characteristics. These business processes were then connected to form three realistically sized business processes. We benchmarked our methods against traditional heuristics and existing resource allocation methods. The results show that our methods learn adaptive resource allocation policies that outperform or are competitive with the benchmarks in five out of six individual business processes. The DRL approach outperforms all benchmarks in all three composite business processes and finds a policy that is, on average, 13.1% better than the best-performing benchmark.
@@ 1 Introduction
In recent years, the main contribution to data-driven optimization of business processes has been seen in Prescriptive Processing Monitoring (PrPM) kubrak_prescriptive_2021 . This research area has seen a strong increase in attention and aims to make recommendations for a case (i.e., an instance of the process) that lead to the best outcome. For example, several PrPM methods recommend the best resource to perform a specific activity for running cases to minimize the case’s completion time sindhgatta_context-aware_2016 ; wibisono_–fly_2015 or cycle time thomas_online_2017 . However, the main limitation of PrPM is that it only optimizes for individual cases and does not consider the effect of intervention for one case on other cases kubrak_prescriptive_2021 .
Data-driven business process optimization (BPO) methods aim to optimize the performance of the complete process and not just individual cases (e.g.,  park_prediction-based_2019 ; huang_reinforcement_2011 ; firouzian_cycle_2019 ; zbikowski_deep_2023 ). In BPO, operational decisions are made to optimize process key performance indicator(s) of interest, such as the mean cycle time or throughput, instead of single case outcome. Take an example of a business process given in Fig. 1. This process has two activities, Accept Application (AA) and Reject Application (RA), and two resources, $r_{1}$ and $r_{2}$. Activity AA can only be performed by resource $r_{1}$, and the activity RA can be performed by both resources. In a situation where two cases are present, case A at activity AA and case B at activity RA, a PrPM method would recommend using resource $r_{1}$ for both cases as it the most efficient resource. However, assigning resource $r_{1}$ to case B means that case A needs to wait until resource $r_{1}$ becomes available again. From a process perspective, assigning resource $r_{2}$ to case B and resource $r_{1}$ to case A would be efficient as both cases can be processed simultaneously and are not competing with each other over the available resources.
@@
The online resource allocation problem in business processes is a dynamic task assignment problem spivey_dynamic_2004 . Several approaches have been proposed to solve this problem. Park and Song park_prediction-based_2019 predicted the next activities and processing times of activity-resource pairs in running cases and solved a min-cost max-flow network problem to find optimal resource assignments. Huang et al. huang_reinforcement_2011 used Reinforcement Learning (RL) to learn the effective resource assignments. This approach was extended by Firouzian et al. firouzian_cycle_2019 by including entropy-based features mined from the event log. However, the tabular Q-learning method used in these papers is unsuitable for problems with large state spaces, such as business processes containing numerous activities and resources. Žbiwkowski et al. zbikowski_deep_2023 partly solved this problem using Deep Reinforcement Learning (DRL), where a neural network replaces the Q-table to maximize the number of completed cases. Their method, however, learns feasible actions through experience, which becomes computationally expensive in larger business processes and may not converge to competitive policies.
In this paper, we present two learning methods for resource allocation in business processes to minimize the mean cycle time of cases. The first method leverages DRL to learn the goodness of assignments by interacting with the simulator of a business process. It can autonomously derive high-dimensional features directly from the state information. The second method is a score-based value function approximation (SVFA) method, which leverages domain expertise to curate process features and learns weights for these features to rank different resource assignments.
To evaluate our models, we first created six distinct business process models with archetypal business process flows and characteristics, such as parallel and choice patterns, varying resource utilization rates, and resource competition. Subsequently, these six business processes were combined in sequential and parallel order to create three realistically-sized business processes. We benchmarked our methods with traditional heuristics and existing BPO methods. The results show that our learning methods outperform or are competitive with the benchmarks in five out of the six business processes. Our DRL method outperforms all of the benchmarks in all three of the composite business processes.
@@ 2 Related work
In business process management, much research has been done on the behavior and allocation of resources. Several methods study the behavior of resources resulting from the current resource allocation policy. This behavior can be evaluated using a set of resource behavior indicators, such as resource preference and productivity pika_extensible_2014 , or from different perspectives huang_resource_2012 . López-Pintado et al. lopez-pintado_prosimos_2023 developed specialized tools for discovering resource profiles to differentiate resource behavior within roles. Furthermore, the current resource allocation policy may also be used to assess the efficiency of a business process duran_analysis_2019 .
@@
Several rule-based methods use process mining techniques to determine static resource allocation strategies. Xu et al. xu_resource_2008 assigned resources to activities based on the workflow of the process model. Other methods use resource-related information mined from the event log in various dimensions to create a ranking of resource assignments, which is used to assign resources to activities at run-time zhao_optimization_2015 ; arias_framework_2016 ; kuchar_automatic_2016 . A generic genetic algorithm was also used for a resource assignment scheme for a specific process based on colored Petri nets si_petri_2018 . These methods provide a generalized resource allocation scheme for the business process model but are not adaptive to the specific circumstances of the process, such as the number of available tasks and resources.
@@
Dynamic methods make assignments based on the current circumstances of the process. Park and Song park_prediction-based_2019 proposed a scheduling approach that uses an offline prediction model to predict the next activities and processing times of resource assignments of a running case. These predictions serve as input for the minimum cost and maximum flow algorithm, which finds optimal resource-activity pairs at each decision step. However, the method only optimizes for the current decision step and does not consider the long-term effects of decisions.
@@
In recent years, RL has become increasingly popular and has also been applied to business process optimization. RL is an optimization method for sequential decision-making problems, such as resource allocation, where a decision-making agent learns a policy by taking actions in an environment sutton_reinforcement_2018 . At each decision step, the agent receives specific information from the business process called the state and, based on the action taken, receives a reward signal. Huang et al. huang_reinforcement_2011 proposed RLRAM, a method that uses the RL algorithm Q-learning, where a Q-table stores the expected future rewards for each state-action pair. They use action-dependent features in their state representation and train their agent to minimize the flow time of the process. This method was extended by Firouzian et al. firouzian_cycle_2019 , who added entropy-based features mined from the event log. While this method works relatively well for problems with small state spaces, the method becomes unsuitable for processes with large state spaces, as separate Q-values are stored for each state. Żibowski et al. zbikowski_deep_2023 solved this issue by DRL to maximize the number of completed cases in a business process. The DRL approach replaces the Q-table with a neural network approximating the value function. However, their method must learn which actions are feasible in each state, which is computationally expensive and may not converge to competitive policies.
@@
Manufacturing processes are closely related to business processes, and numerous methods have been proposed for resource allocation in manufacturing. Many heuristics exist, such as Earliest Due Date or First-In-First-Out sels_comparison_2012 . While these heuristics are robust, explainable, and computationally efficient, they are less effective in situations for which they are not explicitly developed nie_gep-based_2013 . DRL has also been applied in job shop scheduling cunha_deep_2020 . For example, Luo luo_dynamic_2020 trained an agent to select dispatching rules, and Song et al. song_flexible_2023 solved a dynamic assignment problem by employing graph neural networks and DRL to capture complex structural relationships between different resources and jobs. However, these methods often assume the sequence of activities in a case is known, which is not true for business processes.
@@
In this paper, we present two dynamic methods that address the limitations of existing methods for resource allocation in business processes. Our methods are adaptive to the current circumstances of the process and can learn near-optimal policies for realistically sized business processes.
@@ 3 Preliminaries
This section introduces definitions and concepts which we will use in this paper. Section 3.1 provides definitions related to resource allocation in business process, and Section 3.2 formalizes the business process model’s execution.
@@ 3.1 Business process models
This paper considers business process models describing the activities that can be performed in an organization and the resources that can perform them, including the arrival process of new cases and the processing times of activities (e.g., the business process in Fig. 1). We assume the process model can be simulated or executed according to some execution semantics, which makes the business process observable. Specifically, at any point in time, we can observe the execution state of the business process and events that change the state of the process, such as the completion of activities or the arrival of new cases. We formalize these aspects using the following definitions.
Definition 1 (Event, Trace)
Let $\mathcal{A}$ be the set of activities, $\mathcal{C}$ the set of cases, $\mathcal{R}$ the set of resources, and $\mathcal{T}$ the time domain. An event $e=(a,c,t,r,l)$ is mapped to an activity $a\in\mathcal{A}$, has a unique case identifier $c\in\mathcal{C}$, is executed at timestamp $t\in\mathcal{T}$ by resource $r\in\mathcal{R}$, and has a life cycle stage $l\in\{start,complete\}$ to indicate if the event represents the start or completion of executing an activity. In line with van der Aalst van_der_aalst_process_2016 , we refer to the different elements $a$, $c$, $t$, $r$ and $l$ of an event $e$ as #$\#_{a}(e)$, #$\#_{c}(e)$, #$\#_{t}(e)$, #$\#_{r}(e)$ and #$\#_{l}(e)$. A trace is a finite, non-empty sequence of events that describes everything that happened in a business case. We define a trace as $\sigma=\langle e_{1},e_{2},\dots,e_{n}\rangle$, where $n$ is the length of the trace $|\sigma|$. Events within a trace are sequential and should be non-decreasing in time such that $\forall_{1\leq i,j\leq n}\#_{t}(e_{i})\geq\#_{t}(e_{j})$.
The process model in Fig. 1 consists of two activities, which can be executed once for each case in the process. We distinguish executions of the same activity, both within and without the same case, using the term activity instance.
@@
An activity instance is an occurrence of an activity that belongs to a specific case. One case may have multiple instances of the same activity during its execution. We define the set of activity instances as $\mathcal{K}$, and each activity instance $k\in\mathcal{K}$ belongs to a case $c$. We define $\mathcal{K}_{c}$ as the activity instances of case $c\in\mathcal{C}$ and $\mathcal{K}_{a}$ as the activity instances of activity $a\in\mathcal{A}$. An activity instance has a life cycle $p\in\{waiting,processing,complete\}$ to indicate whether the activity instance is waiting in the queue, being processed, or is completed. We refer to the life cycle stage of an activity instance as #$\#_{p}(k)$.
To model the execution state of a business process for the purpose of learning what the best resource allocation is in a particular state, we must be able to observe what activity instances are waiting for processing and what resources are available to be assigned to these activity instances.
@@
An unassigned activity instance is an activity waiting in the queue of an activity to be executed by a resource. We define the set of unassigned activity instances as $K=\{k\in\mathcal{K}\mid\#_{p}(k)=waiting\}$. We use subscripts to indicate to what case $K_{c}$ and activity $K_{a}$ an unassigned activity instance belongs.
@@
An available resource is currently not busy executing an activity instance and can be assigned. We define the set of available resources as $R^{+}\subseteq\mathcal{R}$ and the set of unavailable resources as $R^{-}\subseteq\mathcal{R}$ such that $R^{+}\cup R^{-}=\mathcal{R}$ and $R^{+}\cap R^{-}=\emptyset$. A resource is removed from $R^{+}$ and added to $R^{-}$ when executing an activity instance. An unavailable resource is moved from $R^{-}$ to $R^{+}$ after it finishes processing the activity instance to which it was assigned.
The business process execution state contains all information about the condition of the process at a given point in time, such as what resources and activity instances are available, as well as what is currently being processed. From a business process execution state, the process may transition into another state after some time $t$ has elapsed as a consequence of an event happening. These events can be the arrival of a new case, the assignment of an activity instance, the completion of an activity instance, or the completion of a case. Events happen according to the process model. An activity can only be executed if assigned to a resource.
@@
A business process execution state is the set of active cases $C\subseteq\mathcal{C}$, unassigned activity instances $K\subseteq\mathcal{K}$, available and unavailable resources $R^{+},R^{-}\subseteq\mathcal{R}$, and current assignments $B\subseteq K\times R^{-}$ that exist at a particular moment in time $t$.
3.2 Business process execution
The resource allocation problem we aim to solve in this paper is a dynamic task assignment problem spivey_dynamic_2004 . Given a business process execution state, we aim to assign a resource to an unassigned activity instance to optimize some KPI under specific constraints. We define the specific parts of the optimization problem, which are the decision variables, objective, and constraints, in the context of the dynamic task assignment below.
One explicit constraint in a business process is resource eligibility, which says that not every resource can perform each activity due to, for example, a lack of a specific skill set or level of authorization.
Definition 6 (Resource Eligibility)
The process model defines what resources are eligible to perform what activity. We refer to the set of resources that can execute activity $a\in\mathcal{A}$ as $\mathcal{R}_{a}\subseteq\mathcal{R}$, which implies that $r\in\mathcal{R}_{a}$ can also execute each unassigned activity instance $k\in\mathcal{K}_{a}$.
The process model in Fig. 1 defines that both resources can perform activity RA, while only resource $r_{1}$ can perform activity AA. Therefore, the eligible resource set for activity RA is $\mathcal{R}_{RA}=\{r_{1},r_{2}\}$, and for activity AA is $\mathcal{R}_{AA}=\{r_{1}\}$.
The decision variables are the assignments, which are the allocation of an available resource to an unassigned activity instance. The set of assignments only includes assignments that satisfy the resource eligibility constraint (Def. 6). However, not all assignments are possible in any given business process execution state (Def. 5) because there must be an unassigned activity instance (Def. 3) and an available resource to perform it (Def. 4). Under these constraints, we define the set of possible assignments as a subset of the set of assignments.
Definition 7 (Assignment, Possible Assignment)
An assignment is a tuple $(r,k)$ of a resource $r\in\mathcal{R}$ and an activity instance $k\in\mathcal{K}$. Considering resource eligibility, we define the set of assignments as $\mathcal{D}=\{(r,k)\mid\forall a\in\mathcal{A},r\in\mathcal{R}_{a},k\in\mathcal{K}_{a}\}$. However, not every assignment is possible in all business process execution states, as there must be unassigned activity instances, and resources must be available and eligible to perform them. We define the set of possible assignments in a specific business process execution state as $D=\{(r,k)\mid(r,k)\in\mathcal{D},r\in R^{+},k\in K\}$.
After making an assignment, completing an activity instance, or upon the arrival of a new case, the business process execution state (Def. 5) transitions to a new state.
@@ 3.2 Business process execution
A state transition happens each time a decision is made to assign resource $r$ to unassigned activity instance $k$. First, the business process execution state (Def. 5) is adapted by changing the set of unassigned activity instances to $K-\{k\}$, the set of available resources to $R^{+}-\{r\}$, the set of unavailable resources to $R^{-}\cup\{r\}$, and the set of current assignments to $B\cup\{(r,k)\}$. The set of possible assignments, $D$, also changes depending on $R^{+}$ and $K$. The state also changes after an activity instance completes, which releases the resource of assignment $(r,k)$ such that $R^{+}\cup\{r\}$, $R^{-}-\{r\}$ and $B-\{(r,k)\}$. Furthermore, if after completing an activity instance $k\in\mathcal{K}_{c}$, case $c$ is not completed, a new unassigned activity instance $k$ is generated, i.e., $K\cup\{k\}$. Upon a new case arrival, an unassigned activity instance is also generated.

### v3

@@ Abstract
Efficient allocation of resources to activities is pivotal in executing business processes but remains challenging. While resource allocation methodologies are well-established in domains like manufacturing, their application within business process management remains limited. Existing methods often do not scale well to large processes with numerous activities or optimize across multiple cases. This paper aims to address this gap by proposing two learning-based methods for resource allocation in business processes to minimize the average cycle time of cases. The first method leverages Deep Reinforcement Learning (DRL) to learn policies by allocating resources to activities. The second method is a score-based value function approximation approach, which learns the weights of a set of curated features to prioritize resource assignments. We evaluated the proposed approaches on six distinct business processes with archetypal process flows, referred to as scenarios, and three realistically sized business processes, referred to as composite business processes, which are a combination of the scenarios. We benchmarked our methods against traditional heuristics and existing resource allocation methods. The results show that our methods learn adaptive resource allocation policies that outperform or are competitive with the benchmarks in five out of six scenarios. The DRL approach outperforms all benchmarks in all three composite business processes and finds a policy that is, on average, 12.7% better than the best-performing benchmark.
@@ 1 Introduction
In recent years, the main contribution to data-driven optimization of business processes has been seen in Prescriptive Processing Monitoring (PrPM) [1]. This research area has seen a strong increase in attention and aims to make recommendations for a case (i.e., an instance of the process) that lead to the best outcome. For example, several PrPM methods recommend the best resource to perform a specific activity for running cases to minimize the case’s completion time [2, 3] or cycle time [4]. However, the main limitation of PrPM is that it only optimizes for individual cases and does not consider the effect of intervention for one case on other cases [1].
Data-driven business process optimization (BPO) methods aim to optimize the performance of the complete process and not just individual cases (e.g.,  [5, 6, 7, 8]). In BPO, operational decisions are made to optimize process key performance indicator(s) of interest, such as the mean cycle time or throughput, instead of single case outcome. Take as an example the business process in Figure 1. This process has two activities, Accept Application (AA) and Reject Application (RA), and two resources, $r_{1}$ and $r_{2}$. Activity AA can only be performed by resource $r_{1}$, and the activity RA can be performed by both resources. In a situation where two cases are present, case A at activity AA and case B at activity RA, a PrPM method would recommend using resource $r_{1}$ for both cases as it the most efficient resource. However, assigning resource $r_{1}$ to case B means that case A needs to wait until resource $r_{1}$ becomes available again. From a process perspective, assigning resource $r_{2}$ to case B and resource $r_{1}$ to case A would be efficient as both cases can be processed simultaneously and are not competing with each other over the available resources.
@@
The online resource allocation problem in business processes is a dynamic task assignment problem [9]. Several approaches have been proposed to solve this problem. Park and Song [5] predicted the next activities and processing times of activity-resource pairs in running cases and solved a min-cost max-flow network problem to find optimal resource assignments. Huang et al. [6] used Reinforcement Learning (RL) to learn the effective resource assignments. This approach was extended by Firouzian et al. [7] by including entropy-based features mined from the event log. However, the tabular Q-learning method used in these papers is unsuitable for problems with large state spaces, such as business processes containing numerous activities and resources. Žbiwkowski et al. [8] partly solved this problem using Deep Reinforcement Learning (DRL), where a neural network replaces the Q-table to maximize the number of completed cases. Their method, however, learns feasible actions through experience, which becomes computationally expensive in larger business processes and may not converge to competitive policies.
In this paper, we present two learning methods for resource allocation in business processes to minimize the mean cycle time of cases. The first method leverages DRL to learn effective policies by interacting with the simulator of a business process. It can autonomously derive high-dimensional features directly from the state information. The second method is a score-based value function approximation (SVFA) method, which leverages domain expertise to curate process features and learns weights for these features to rank different resource assignments.
To evaluate our models, we first created six distinct business process models, referred to as scenarios, with archetypal business process flows and characteristics, such as parallel and choice patterns, varying resource utilization rates, and resource competition. Subsequently, these scenarios were combined in sequence, in sequence with reversed order, and in parallel to create three realistically sized business processes. We benchmarked our methods with traditional heuristics and existing BPO methods [5, 8]. The results show that our learning methods outperform or are competitive with the benchmarks in five out of the six business processes. Our DRL method outperforms all of the benchmarks in all three of the composite business processes.
@@ 2 Related work
In business process management, much research has been done on the behavior and allocation of resources. Several methods study the behavior of resources resulting from the current resource allocation policy. This behavior can be evaluated using a set of resource behavior indicators, such as resource preference and productivity [10], or from different perspectives [11]. López-Pintado et al. [12] developed specialized tools for discovering resource profiles to differentiate resource behavior within roles. Furthermore, the current resource allocation policy may also be used to assess the efficiency of a business process [13].
@@
Several rule-based methods use process mining techniques to determine static resource allocation strategies. Xu et al. [14] assigned resources to activities based on the workflow of the process model. Other methods use resource-related information mined from the event log in various dimensions to create a ranking of resource assignments, which is used to assign resources to activities at run-time [15, 16, 17]. A generic genetic algorithm was also used for a resource assignment scheme for a specific process based on colored Petri nets [18]. These methods provide a generalized resource allocation scheme for the business process model but are not adaptive to the specific circumstances of the process, such as the number of available tasks and resources.
@@
Dynamic methods make assignments based on the current circumstances of the process. Park and Song [5] proposed a scheduling approach that uses an offline prediction model to predict the next activities and processing times of resource assignments of a running case. These predictions serve as input for the minimum cost and maximum flow algorithm, which finds optimal resource-activity pairs at each decision step. However, the method only optimizes for the current decision step and does not consider the long-term effects of decisions.
@@
In recent years, RL has become increasingly popular and has also been applied to business process optimization. RL is an optimization method for sequential decision-making problems, such as resource allocation, where a decision-making agent learns a policy by taking actions in an environment [19]. At each decision step, the agent receives specific information from the business process called the state and, based on the action taken, receives a reward signal. Huang et al. [6] proposed RLRAM, a method that uses the RL algorithm Q-learning, where a Q-table stores the expected future rewards for each state-action pair. They use action-dependent features in their state representation and train their agent to minimize the flow time of the process. This method was extended by Firouzian et al. [7], who added entropy-based features mined from the event log. While this method works relatively well for problems with small state spaces, the method becomes unsuitable for processes with large state spaces, as separate Q-values are stored for each state. Żibowski et al. [8] solved this issue by DRL to maximize the number of completed cases in a business process. The DRL approach replaces the Q-table with a neural network approximating the value function. However, their method must learn which actions are feasible in each state, which is computationally expensive and may not converge to competitive policies.
@@
Manufacturing processes are closely related to business processes, and numerous methods have been proposed for resource allocation in manufacturing. Many heuristics exist, such as Earliest Due Date or First-In-First-Out [20]. While these heuristics are robust, explainable, and computationally efficient, they are less effective in situations for which they are not explicitly developed [21]. DRL has also been applied in job shop scheduling [22]. For example, Luo [23] trained an agent to select dispatching rules, and Song et al. [24] solved a dynamic assignment problem by employing graph neural networks and DRL to capture complex structural relationships between different resources and jobs. However, these methods often assume the sequence of activities in a case is known, which is not true for business processes.
@@
In this paper, we present two dynamic methods that address the limitations of existing methods for resource allocation in business processes. Our DRL method extends current research on DRL [8] by improving scalability and efficiency while optimizing a more complex objective. Our SVFA approach, which employs a ranking of assignments similar to the methods in [15, 16, 17], improves these techniques by learning optimal feature weights and dynamically adjusting the ranking to the state of the business process. Our methods are adaptive to the current circumstances of the process and can learn near-optimal policies for realistically sized business processes.
@@ 3 Preliminaries
This section introduces definitions and concepts which we will use in this paper. Section 3.1 provides essential definitions for understanding business process models, Section 3.2 defines how business processes are executed based on their model, and Section 3.3 formalizes the problem of allocating resources to activities in a business process.
@@ 3.1 Business process models
This paper considers business process models describing the activities that can be performed in an organization and the resources that can perform them, including the arrival process of new cases and the processing times of activities (e.g., the business process in Figure 1). We assume the process model can be simulated or executed according to some execution semantics, which makes the business process observable. Specifically, at any point in time, we can observe the execution state of the business process and events that change the state of the process, such as the completion of activities or the arrival of new cases. We formalize these aspects using the following definitions.
Let $\mathcal{A}$ be the set of activities in a business process, $\mathcal{C}$ the set of cases, $\mathcal{T}$ the time domain, $\mathcal{R}$ the set of resources, and $\{start,complete\}$ be the set of activity life cycle transitions that indicate starting and completing an activity.
Definition 1 (Event)
$\mathcal{E}=\mathcal{A}\times\mathcal{C}\times\mathcal{T}\times\mathcal{R}\times\{start,complete\}$ is the set of all possible events. An event is a tuple $(a,c,t,r,l)\in\mathcal{E}$.
In line with van der Aalst [25], we refer to the different elements (activity, case, time, resource, and life cycle transition) of an event $e$ as #$\#_{a}(e)$, #$\#_{c}(e)$, #$\#_{t}(e)$, #$\#_{r}(e)$ and #$\#_{l}(e)$. A trace is a finite, non-empty sequence of events that describes everything that happened in a business case.
Definition 2 (Trace)
For $e_{1},e_{2},\ldots,e_{n}\in\mathcal{E}$, $\langle e_{1},e_{2},\dots,e_{n}\rangle$ is a trace with length $n$. Events in a trace must be non-decreasing in time: $\forall_{1\leq i\leq j\leq n}\#_{t}(e_{i})\leq\#_{t}(e_{j})$. The length of a trace $\sigma$ is also denoted as $|\sigma|$.
The process model in Figure 1 consists of two activities. Activities can be executed multiple times. We refer to the occurrence of an activity as an activity instance. Note that, while in the example of Figure 1 an activity is executed only once for each case, activities can be executed multiple times for the same case.
@@
Let $\mathcal{K}$ be the set of all activity instances. $\mathcal{K}_{c}$ is the set of all activity instances of case $c\in\mathcal{C}$ and $\mathcal{K}_{a}$ is set of all activity instances of activity $a\in\mathcal{A}$. An activity instance $k\in\mathcal{K}$ has a life cycle state #$\#_{p}(k)\in\{waiting,processing,complete\}$.
We also use subscripts to indicate to what case $K_{c}\in\mathcal{K}_{c}$ and activity $K_{a}\in\mathcal{K}_{a}$ an activity instance belongs.
To model the execution state of a business process for the purpose of learning what the best resource allocation is in a particular state, we must be able to observe what activity instances are waiting for processing and what resources are available to be assigned to these activity instances. An unassigned activity instance is an activity waiting in the queue of an activity to be executed by a resource.
@@
An unassigned activity instance is an element from the set $\{k\in\mathcal{K}\mid\#_{p}(k)=waiting\}$.
An available resource is a resource that is currently not executing an activity instance and therefore can be assigned to execute one.
@@
$R^{+}\subseteq\mathcal{R}$ is the set of available resources and $R^{-}\subseteq\mathcal{R}$ the set of unavailable resources, such that $R^{+}\cup R^{-}=\mathcal{R}$ and $R^{+}\cap R^{-}=\emptyset$.
A resource is removed from $R^{+}$ and added to $R^{-}$ when it starts executing an activity instance. A resource is moved from $R^{-}$ to $R^{+}$ when it finishes processing the activity instance to which it was assigned.
A business process defines which resource is eligible to perform which activity. Not every resource can perform each activity due to, for example, a lack of a specific skill set or level of authorization.
Definition 6 (Resource Eligibility)
For an activity $a\in\mathcal{A}$, $\mathcal{R}_{a}\subseteq\mathcal{R}$ is the set of resources that can execute $a$, which implies that $r\in\mathcal{R}_{a}$ can also execute each $k\in\mathcal{K}_{a}$.
For example, the process model from Figure 1 defines that both resources can perform activity RA, while only resource $r_{1}$ can perform activity AA. Therefore, the eligible resource set for activity RA is $\mathcal{R}_{RA}=\{r_{1},r_{2}\}$, and for activity AA is $\mathcal{R}_{AA}=\{r_{1}\}$.
3.2 Business Process Execution
The way in which business processes can be executed is typically defined in terms of how the state of a business process can transition, depending on what happens in the business process. The way in which business processes can transition in state has been well-defined for different kinds of business process modeling notations, such as BPMN and Petri nets. Accordingly, we assume that a business process model exists that determines when a new case can arrive and which activity (instance) can be executed when a new case arrives or when the execution of another activity instance completes. We extend this with the effect that resource assignment has on state transitions.
The business process execution state contains all information about the condition of the process at a given point in time, such as what resources and activity instances are available, as well as what is currently being processed. From a business process execution state, the process may transition into another state after some time $t$ has elapsed as the consequence of an event occurring. These events can be the arrival of a new case, the assignment of an activity instance, the completion of an activity instance, or the completion of a case. Events happen according to the process model. An activity can only be executed if assigned to a resource.
@@ 3.2 Business Process Execution
Let $C\subseteq\mathcal{C}$ be a set of cases, $K\subseteq\mathcal{K}$ be a set of unassigned activity instances, $R^{+},R^{-}\subseteq\mathcal{R}$ be sets of assigned and unassigned resources, $B\subseteq Z\times R^{-}$, where $Z=\{k\in\mathcal{K}\mid\#_{p}(k)=processing\}$, be a current assignment of resources to activity instances, and $t\in\mathcal{T}$ a moment in time, the execution state of a business process is a tuple $(C,K,R^{+},R^{-},B,t)$ consisting of those elements. $\mathcal{S}$ is the set of all business process execution states.
In a particular execution state of a business process it is possible to assign resources to activity instances, if they are available and eligible to execute those activity instances.
Definition 8 (Assignment)
An assignment is a tuple $(r,k)$ of a resource $r\in\mathcal{R}$ and an activity instance $k\in\mathcal{K}$. Considering resource eligibility, the set of eligible assignments $\mathcal{D}=\{(r,k)\mid\forall a\in\mathcal{A},r\in\mathcal{R}_{a},k\in\mathcal{K}_{a}\}$. The set of possible assignments in a specific business process execution state as $D=\{(r,k)\mid(r,k)\in\mathcal{D},r\in R^{+},k\in K\}$.
When an assignment is made, the process transitions into a new execution state. This transition occurs due to three things happening: (1) the assignment itself; (2) the completion of the execution of an activity instance; or (3) the arrival of a new case. For (2) and (3), we assume that a business process model exists that describes when cases arrive and which activities can be executed when a case arrives or when another activity completes.
@@
From a business process execution state $(C,K,R^{+},R^{-},B,t)\in\mathcal{S}$, a business process can transition into a new state due to three possible events:
  1. 1.
    By assigning $(r,k)\in D$.
  2. 2.
    By the completion of an activity instance $k\in K$.
  3. 3.
    By the arrival of a new case $c\notin C$.
When assigning $(r,k)\in D$, the business process transitions into a state $(C,K^{\prime},R^{+\prime},R^{-\prime},B^{\prime},t)$, where $K^{\prime}=K-\{k\}$, $R^{+\prime}=R^{+}-\{r\}$, $R^{-\prime}=R^{-}\cup\{r\}$, and $B^{\prime}=B\cup\{(r,k)\}$. When $k\in K$ completes with $(r,k)\in B$, the business process transitions into a state $(C,K^{\prime},R^{+\prime},R^{-\prime},B^{\prime},t)$, where $K^{\prime}=K-\{k\}$, $R^{+\prime}=R^{+}\cup\{r\}$, $R^{-\prime}=R^{-}-\{r\}$ and $B^{\prime}=B-\{(r,k)\}$. Subsequently, according to the business process model, new activity instances may be generated and added to $K^{\prime}$. At any moment, according to the business process model, new cases may arrive, which trigger the addition of a case to the set of cases $C$ and initial activity instances to the set of activity instances $K$.
Definition 10 (State Transition Function)
$\tau:\mathcal{S}\times D\rightarrow\mathcal{S}$ is a transition function that, for a state $S$ and an assignment $(r,k)$ of a resource $r$ to an activity instance $k$ returns the next state $S^{\prime}$ that is derived through Def. 9 and simulation of the business process model.
3.3 Resource allocation problem
The resource allocation problem we aim to solve in this paper is a dynamic task assignment problem [9]. The solution to such a problem is a policy that, given a business process execution state, assigns a resource to an unassigned activity instance.
We can now define a resource allocation policy as follows.
Definition 11 (Resource Allocation Policy)
A resource allocation policy is a function $\pi:\mathcal{S}\rightarrow\mathcal{R}\times\mathcal{A}$, which for a state $(C,K,R^{+},R^{-},B,t)\in\mathcal{S}$ returns a possible assignment $(r,k)\in R^{+}\times K$.
The goal is to find an optimal policy. This is the policy that, when applied to the simulator for a particular duration, optimizes a specific objective. The duration can be flexibly chosen, for example, as a particular duration of simulation time or a particular number of cases. We define the duration by setting a function $done:\mathcal{S}\rightarrow Bool$ that returns true if and only if a business process execution state is a final state.
Definition 12 (Rollout)
Let $S$ be the initial state of a business process, $\pi$ be a policy, $\tau$ be a transition function, and $done$ be the simulation duration function. A rollout $\hat{\sigma}$ is a trace $\sigma$ that is generated by applying $S^{\prime}=\tau(S,\pi(S))$ until $done(S)$.
It is important to see that due to randomness in the business process simulation, for example, in the arrivals and processing times, applying the same policy to the same start state may lead to different rollouts. We define the function that determines the value of a trace.
Definition 13 (Reward, Value)
Let $\hat{\sigma}$ be a rollout. A reward function is a function $r:\mathcal{E}\rightarrow\mathbb{R}$ that assigns a reward to each event that happens in the rollout. The value of a rollout $\hat{\sigma}$ is
$V(\hat{\sigma})=\sum_{i=1}^{|\hat{\sigma}|}r(e)$
Let $\Sigma$ be a multiset of rollouts that is generated by simulation. The expected value of the policy $\pi$ that generated the rollouts
$\hat{V}_{\pi}(\Sigma)=\frac{1}{|\Sigma|}\sum_{\hat{\sigma}\in\Sigma}V(\sigma)$